OpenStudy (anonymous):

The greatest value of \[\frac{4}{4x^2+4x+9}\] is- Ans.1/2 method please!!!!

OpenStudy (anonymous):

firstly complete the square for 4x^2+4x+9

OpenStudy (anonymous):

??

OpenStudy (anonymous):

\[\large 4x^2+4x+9=(2x+1)^2+8\] am i right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

well now use this \[(2x+1)^2+8\ge 8 \\ since\\ (2x+1)^2\ge0\] for any real number x

OpenStudy (anonymous):

what about the numerator?

OpenStudy (anonymous):

\[\huge a>b \\ then \\ \huge \frac{1}{a}<\frac{1}{b}\]

OpenStudy (anonymous):

what will u get?

OpenStudy (anonymous):

i don't know , you solve ahead.

OpenStudy (anonymous):

\[(2x+1)^2+8\ge8 \ \ then \ \ \frac{1}{(2x+1)^2+8} \le \frac{1}{8}\]

OpenStudy (anonymous):

multiply it by 4 and u will get the answer

OpenStudy (anonymous):

A fraction of constant numerator is maximum when the denominator is minimum. As @mukushla is trying to explain to you that the denominator is minimum when x=-1/2. You can verify that the fraction is equal to 1/2 when x=-1/2

OpenStudy (anonymous):

Also you can do this problem using derivatives \[ f(x)=\frac{4}{4 x^2+4 x+9}\\ f'(x)=-\frac{16 (2 x+1)}{\left(4 x^2+4 x+9\right)^2} \] You can deduce that f has a maximum at x =-1/2